40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8557     Entry time: Thu May 9 02:19:53 2013
Author: Jenne 
Type: Update 
Category: Locking 
Subject: 50% BS installed in POP path 

Koji had the good idea of trying to measure the motion of the POP beam, and feeding that signal to PRM yaw to stabilize the motion.  To facilitate this, I have installed a 50% beam splitter before the POP 110/22 PD (so also before the camera). 

Before touching anything, I locked the PRM-ITMY half-cavity so that I had a constant beam at POP.  I measured the POP DC OUT to be 58.16 counts.  I then installed a 1" 50% BS, making sure (using the 'move card in front of optic while watching camera' technique) that I was not close to clipping on the new BS.  I then remeasured POP DC OUT, and found it to be 30.63.  I closed the PSL shutter to get the dark value, which was -0.30 .  This means that I now have a factor of 0.53 less light on the POP110/22 PD.  To compensate for this, I changed the values of the power normalization matrix from 0.01 (MICH) to 0.0189, and 100 (PRCL) to 189.

After doing this, I restored the ITMX and am able to get several tens of seconds of PRMI lock (using AS55Q and REFL33I). 

I found several QPDs in the PD cabinet down the Y arm, but no readout electronics.  The QPD I found is D990272.  I don't really want to spend any significant amount of time hacking something for this together, if Valera can provide a QPD with BNC outputs. For now, I have not installed any DC PD or razor blade (which can be a temporary proxy for a QPD, enough to get us yaw information).

 

ELOG V3.1.3-